Federal Income Tax Withholding

Bob Dunn's weekly gross earnings for the present week were $2,950. Dunn has two exemptions. Using the wage bracket withholding table in Exhibit 2 with a $75 standard withholding allowance for each exemption, what is Dunn's federal income tax withholding? Round your answer to two decimal places.

$

Employee Net Pay

Maxwell Friedman's weekly gross earnings for the week ended April 22 were $1,302, and his federal income tax withholding was $234.36. Assuming the social security rate is 6% and Medicare is 1.5%, what is Maxwell's net pay?

Note: If required, round your answer to two decimal places.

$
Answer #1

Answer 1

Total wage payment 2950

Total allowance (75*2) 150

Amount subject to withholding 2800

Initial withholding = 302.95

+ additional withholding : 28% withholding in excess of 1533 = (28%*(2800-1533)) = 354.76

Federal income tax withholding =657.71

Answer 2

Gross earnings = 1302

Less: Federal income tax withholding = 234.36

Less: Social security (1302*6%) = 78.12

Less: medicare (1302*1.5%) = 19.53

Net pay = 1634.01